# Relational Graph Transformer

### v1

@@ A.2 Node initialization for Subgraph GNN PE in RelGT
The randomly initialized node features ($Z_{\text{random}}$) provide enhanced properties as discussed in Section 3.1. We investigate the alternative approach of using Laplacian PE ($Z_{\text{LapPE}}$) computed over the subgraph instead of random initialization and report these results in Table 4. For these results, we utilized a positional encoding dimension size of $4$. Our findings indicate that $Z_{\text{LapPE}}$ consistently underperforms compared to $Z_{\text{random}}$, while also introducing additional computational overhead ranging from 1.02$\times$ to 3.38$\times$ across the 8 selected tasks in our study. This shows the challenges of using existing PEs such as Laplacian PE in relational entity graphs and signify the use of GNN PE as part of RelGT’s tokenization strategy.
@@
Dataset Task (# train) MAE $\downarrow$ $Z_{\text{random}}$ $Z_{\text{LapPE}}$ Rel Drop $Z_{\text{random}}$ $Z_{\text{LapPE}}$ Runtime Ratio
@@
Dataset Task (# train) AUC $\uparrow$ $Z_{\text{random}}$ $Z_{\text{LapPE}}$ Rel Drop $Z_{\text{random}}$ $Z_{\text{LapPE}}$ Runtime Ratio
@@
A.3 HGT Baseline
In the main experiments (Section 4), we use the Heterogeneous Graph Transformer (HGT) [20] as a graph transformer (GT) baseline, and report results for two variants to demonstrate the advantages of RelGT over existing GT models. Specifically, we consider the standard HGT model and an enhanced version, HGT+PE, which incorporates Laplacian positional encodings (LapPE). These positional encodings are computed on sampled subgraphs rather than the full graph.
For implementation, we use the HGTConv layer from PyTorch Geometric [14] and integrate it into the RDL pipeline [42] by replacing the default GNN module. Both variants use 4 attention heads and 2 layers, similar to the configuration of the GNN module in RDL, with residual connections and layer normalization applied between layers. For the HGT+PE variant, we use LapPE of dimension 4 for all tasks, except for rel-amazon item-ltv and rel-hm item-sales, where we use dimension 2. Notably, because the relational entity graphs are heterogeneous, the Laplacian positional encodings is computed multiple times for each node type, unlike the original homogeneous setting for which LapPE was designed [10].
@@ A.3 HGT Baseline
rel-event user-attendance $-2.85$ 4.36 37.57
rel-trial study-adverse $-2.03$ 9.72 15.02
@@
rel-avito user-visits $-1.24$ 17.16 40.07
@@
rel-trial study-outcome $-0.21$ 4.09 4.83
@@
rel-amazon item-churn $-0.20$ 75.51 152.06
@@
Average $-0.05$ 52.28 128.64
A.4 Detailed Results
In Table 6, we report the full results of different configurations we tuned for RelGT, particularly on the smaller datasets with lesser than a million training nodes. Table 7 provides the full scores for the RelGT component study in Table 2, while Table 8 provides the supporting results for Figure 4. Finally, we provide the elaborated version of the Tables 1(a) and 1(b) in Tables 9 and 9, respectively.
A.5 Resource Information.
We implement RelGT using PyTorch framework [39], PyTorch Geometric framework [14] and adapt the codebase of relational deep learning [42] https://github.com/snap-stanford/relbench. All our experiments are conducted on an NVIDIA A100 GPU server with 8 GPU nodes.
@@ A.4 Detailed Results
rel-f1 driver-position Test 4.022 4.1598 4.2358 3.9170 2.61
Val 3.193 3.3517 2.9894 3.3257
rel-avito ad-ctr Test 0.041 0.0441 0.0494 0.0345 15.85
Val 0.037 0.0409 0.0456 0.0314
rel-event user-attendance Test 0.258 0.2635 0.2562 0.2502 2.79
Val 0.255 0.2617 0.2574 0.2548
rel-trial study-adverse Test 44.473 43.3253 42.4622 43.9923 1.08
Val 46.290 45.9957 45.7966 46.2148
site-success Test 0.400 0.4374 0.4431 0.3263 18.43
Val 0.401 0.4198 0.4245 0.3593
rel-amazon user-ltv Test 14.313 15.3804 15.9296 14.2665 0.32
Val 12.132 13.1017 13.5599 12.1151
item-ltv Test 50.053 56.1384 55.6211 48.9222 2.26
Val 45.1401 51.2139 50.3468 43.8161
rel-stack post-votes Test 0.065 0.0679 0.0680 0.0654 -0.62
Val 0.059 0.0617 0.0618 0.0592
rel-hm item-sales Test 0.056 0.0655 0.0641 0.0536 4.29
Val 0.065 0.0749 0.0735 0.0627
@@
rel-f1 driver-dnf Test 0.7262 0.7142 0.7109 0.7587 4.48
Val 0.7136 0.7678 0.7318 0.6804
driver-top3 Test 0.7554 0.6389 0.8340 0.8352 10.56
Val 0.7764 0.6659 0.6079 0.7958
rel-avito user-clicks Test 0.6590 0.6584 0.6387 0.6830 3.64
Val 0.6473 0.5977 0.5656 0.6649
user-visits Test 0.6620 0.6426 0.6507 0.6678 0.88
Val 0.6965 0.6696 0.6732 0.7024
rel-event user-repeat Test 0.7689 0.6717 0.6590 0.7609 -1.04
Val 0.7125 0.6247 0.5974 0.7285
user-ignore Test 0.8162 0.8348 0.8161 0.8157 -0.06
Val 0.9170 0.8896 0.8940 0.8868
rel-trial study-outcome Test 0.6860 0.5679 0.5691 0.6861 0.01
Val 0.6818 0.5985 0.5925 0.6678
rel-amazon user-churn Test 0.7042 0.6608 0.6589 0.7039 -0.04
Val 0.7045 0.6639 0.6622 0.7036
item-churn Test 0.8281 0.7824 0.7840 0.8255 -0.31
Val 0.8239 0.7845 0.7846 0.8220
rel-stack user-engagement Test 0.9021 0.8898 0.8852 0.9053 0.35
Val 0.9059 0.8914 0.8847 0.9033
user-badge Test 0.8986 0.8652 0.8518 0.8632 -3.94
Val 0.8886 0.8760 0.8691 0.8741
rel-hm user-churn Test 0.6988 0.6773 0.6491 0.6927 -0.87
Val 0.7042 0.6814 0.6502 0.6988

### v2

@@ A.2 Node initialization for Subgraph GNN PE in RelGT
[…]
@@
Dataset Task (# train) MAE $\downarrow$ $Z_{\text{random}}$ $Z_{\text{LapPE}}$ % Rel Drop $Z_{\text{random}}$ $Z_{\text{LapPE}}$ Runtime Ratio
@@
Dataset Task (# train) AUC $\uparrow$ $Z_{\text{random}}$ $Z_{\text{LapPE}}$ % Rel Drop $Z_{\text{random}}$ $Z_{\text{LapPE}}$ Runtime Ratio
@@
A.3 Learnable Spatio-Temporal PE
In this section, we explore a learnable spatio-temporal positional encoding (PE) for RelGT. Instead of using the relative time encoder (Eqn. 4), we use the ‘relative time’ term to initialize nodes in the Subgraph GNN PE, where relative time $\tau(v_{j})-\tau(v_{i})$ denotes the temporal difference between neighbor node $v_{j}$ and seed node $v_{i}$. This approach repurposes the Subgraph GNN PE as a learnable spatio-temporal PE, which is defined as:
$h_{\text{stpe}}(v_{j})=\text{GNN}(A_{\text{local}},Z_{\text{relative\_time}})_{j}\in\mathbb{R}^{d}$ (11)
where $z_{j,\text{relative\_time}}=\tau(v_{j})-\tau(v_{i})$. The token representation, then, becomes:
${h}_{\text{token}}(v_{j})=O\cdot\left[{h}_{\text{feat}}(v_{j})\,||\,{h}_{\text{type}}(v_{j})\,||\,{h}_{\text{hop}}(v_{i},v_{j})\,||\,{h}_{\text{stpe}}(v_{j})\right]$ (12)
where $||$ denotes the concatenation of the individual encoder outputs, and $O\in\mathbb{R}^{4d\times d}$ is a learnable matrix to mix the embeddings.
Table 5 presents our evaluation results over three regression and five classification tasks in RelBench. Across all tasks, replacing the original temporal encoder and subgraph GNN PE encoder with the Spatio-Temporal PE leads to a consistent performance decline in RelGT.
Table 5: Performance comparison of RelGT (Full) with the Spatio-Temporal PE (Eqns. 11 - 12). Negative scores suggest performance drop with the spatio-temporal PE in RelGT.
Dataset Task MAE $\downarrow$ RelGT (Full) RelGT (Spatio-Temporal PE) % Rel. Diff
rel-avito ad-ctr Test 0.0345 0.0355 -2.90
Val 0.0314 0.0315
rel-trial site-success Test 0.3262 0.3554 -8.95
Val 0.3593 0.3883
rel-hm item-sales Test 0.0536 0.0630 -17.54
Val 0.0627 0.0718
Dataset Task AUC $\uparrow$ RelGT (Full) RelGT (Spatio-Temporal PE) % Rel. Diff
rel-avito user-clicks Test 0.6830 0.6465 -5.34
Val 0.6649 0.6519
user-visits Test 0.6678 0.6641 -0.55
Val 0.7024 0.7017
rel-event user-ignore Test 0.8157 0.8152 -0.06
Val 0.8868 0.8870
rel-trial study-outcome Test 0.6861 0.6537 -4.72
Val 0.6678 0.6757
rel-amazon user-churn Test 0.7039 0.7036 -0.04
Val 0.7036 0.7037
Table 6: RelGT results on entity classification tasks in RelBench compared with Griffin (Wang et al., 2025). AUC is the performance metric. Higher is better.
Dataset Task Griffin RelGT (ours) % Rel. Gain
rel-f1 driver-dnf 0.745 0.7587 1.84
driver-top3 0.825 0.8352 1.24
rel-avito user-clicks 0.630 0.6830 8.41
user-visits 0.650 0.6678 2.74
rel-trial study-outcome 0.689 0.6861 -0.42
rel-amazon user-churn 0.700 0.7039 0.56
item-churn 0.811 0.8255 1.79
rel-stack user-engagement 0.898 0.9053 0.81
user-badge 0.870 0.8632 -0.78
rel-hm user-churn 0.683 0.6927 1.42
A.4 Comparison with Griffin
In this section, we compare RelGT with Griffin (Wang et al., 2025), which is a GNN based relational foundation model that integrates unified feature encoders, cross-attention and hierarchical message passing to process relational entity graphs from diverse domains. We report the results in Table 6 for 10 entity classification tasks in RelBench, where Griffin is finetuned on each task following the same procedure used to train RelGT. RelGT outperforms Griffin on 8 out of 10 tasks with relative gains of up to 8.41%, demonstrating the advantages of a Graph Transformer backbone for processing relational entity graphs.
A.5 HGT Baseline
In the main experiments (Section 4), we use the Heterogeneous Graph Transformer (HGT) (Hu et al., 2020) as a graph transformer (GT) baseline, and report results for two variants to demonstrate the advantages of RelGT over existing GT models. Specifically, we consider the standard HGT model and an enhanced version, HGT+PE, which incorporates Laplacian positional encodings (LapPE). These positional encodings are computed on sampled subgraphs rather than the full graph.
For implementation, we use the HGTConv layer from PyTorch Geometric (Fey & Lenssen, 2019) and integrate it into the RDL pipeline (Robinson et al., 2024) by replacing the default GNN module. Both variants use 4 attention heads and 2 layers, similar to the configuration of the GNN module in RDL, with residual connections and layer normalization applied between layers. For the HGT+PE variant, we use LapPE of dimension 4 for all tasks, except for rel-amazon item-ltv and rel-hm item-sales, where we use dimension 2. Notably, because the relational entity graphs are heterogeneous, the Laplacian positional encodings is computed multiple times for each node type, unlike the original homogeneous setting for which LapPE was designed (Dwivedi et al., 2020).
@@ A.5 HGT Baseline
rel-event user-attendance 2.85-2.85 4.36 37.57
rel-trial study-adverse 2.03-2.03 9.72 15.02
@@
rel-avito user-visits 1.24-1.24 17.16 40.07
@@
rel-trial study-outcome 0.21-0.21 4.09 4.83
@@
rel-amazon item-churn 0.20-0.20 75.51 152.06
@@
Average 0.05-0.05 52.28 128.64
A.6 Detailed Results
In Table 8, we report the full results of different configurations we tuned for RelGT, particularly on the smaller datasets with lesser than a million training nodes. Table 9 provides the full scores for the RelGT component study in Table 2, while Table 10 provides the supporting results for Figure 4. Finally, we provide the elaborated version of the Tables 1(a) and 1(b) in Tables 11 and 12, respectively.
A.7 Computational Complexity and Resource Information.
RelGT has $O(K^{2}\cdot d)$ complexity for local attention and $O(K\cdot B\cdot d)$ for global attention per node, where $K$ is local context size, $B$ is number of global centroids, and $d$ is hidden dimension. We implement RelGT using PyTorch framework (Paszke, 2019), PyTorch Geometric framework (Fey & Lenssen, 2019) and adapt the codebase of relational deep learning (Robinson et al., 2024) https://github.com/snap-stanford/relbench. All our experiments are conducted on an NVIDIA A100 GPU server with 8 GPU nodes.
@@ A.6 Detailed Results
[…]
@@
[…]